OpenStudy (anonymous):

Which points are on the line given by the equation y = -5x + 3? (-2,7) (2,-7) (-1,8) (4,-23)

OpenStudy (anonymous):

Plug in the values for x and y and see if the set of points matches the equation

OpenStudy (anonymous):

For the first set of points, you would have 7 = -5(-2) + 3 7 = 10 + 13 // not on the line

OpenStudy (anonymous):

y = -5(2) = 3 -7 = -10 + 3 = -7 // on the line can you do the rest?

OpenStudy (anonymous):

@santiej, that help?

OpenStudy (anonymous):

Yes, it does.

OpenStudy (anonymous):

But how do I know if the point that I get is on the line?

OpenStudy (anonymous):

You know by if y satisfies the equation for the points

OpenStudy (anonymous):

That is why we plugged in a set of points and checked to see if it was on the line or not

OpenStudy (anonymous):

If it doesn't satisfy the equation, it is not on the line

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

Not really.

OpenStudy (anonymous):

How do I know if the equation is satisfied?

OpenStudy (anonymous):

So let's take the first point: (-2, 7), when we plug in 7 for y and -2 in for x, if the two equal each other then that point is on the line

OpenStudy (anonymous):

That might help you to understand

OpenStudy (anonymous):

So, if y and x are equal then I know it is on the line?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Ok, thank you.

OpenStudy (anonymous):

because it satisfies the equation

OpenStudy (anonymous):

Thank you for taking the time to help me.

OpenStudy (anonymous):

Np. I'm glad you got it. :)